Please help me to solve it:
Two times a number plus its reciprocal is 129/8.What is the number(s)?

Three times a number minus three times its reciprocal is 148/7.What is the number(s)?
Thank you.

Answer by mananth(16949)

2x + 1/x = 129/8
(2x^2+1)/x=129/8
cross multiply
8(2x^2+1)=129x
16x^2+8=129x
16x^2-129x+8=0
16x^2-128x-x+8=0
16x(x-8)-1(x-8)=0
(x-8)(16x-1)=0
x= 8 OR 1/16

3x-3/x=148/7
(3x^2-3)/x= 148/7
cross multiply
7(3x^2-3)=148x
21x^2-21=148x
21x^2-148x-21=0
solve by quadratic formula